OpenStudy (anonymous):

Cna you give me the steps and answers to -1X^(2) + 7X - 12 = 0

OpenStudy (anonymous):

-x^2 + 7x -12 = 0 (-x+4)(x-3) = 0 *Check answer by multiplying out: -x^2 +3x+4x -12 = 0 Therefore, x=4, 3